The perimeter of my garden is 32 feet.If my garden is 10 feet wide how long is it.
Answer by macston(5194)

.
L=length; W=width=10 feet; P=perimeter
P=2(L+W)
32 feet=2(L+10 feet) Divide each side by 2.
16 feet=L+10 feet Subtract 10 feet from each side.
6 feet=L ANSWER: The length of the garden is 6 feet.